Exercise 16.2  Determining the maturity value of notes.

Compute the maturity value for each of the following notes:

  • A note payable with a face amount of $25,000, dated June 15, 2019, due in three months, bearing interest at 7 percent.
  • A note payable with a face amount of $22,000, dated May 5, 2019, due in 45 days, bearing interest at 8 percent.
Answer #1

Maturity value = Principal + Interest

= 25,000 + (25,000*7%*3/12)

= 25,000 + 437.50

= 25,437.50

Maturity value

= 22,000 + (22,000*8%*45/360)

= 22,000 + 220

= 22,220